OpenStudy (anonymous):

Assume that you built a robot with a 0.0400s reaction time. How far down (in centimeters) would a ruler fall before being caught by the robot?

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

would u use the formula d=1/2at^2+vi(t)

OpenStudy (anonymous):

Certainly! Would would a be, as well as the initial velocity.

OpenStudy (anonymous):

a is 981 and initial velocity is 0?

OpenStudy (anonymous):

Exactly!

OpenStudy (anonymous):

ok one more question if u have time

OpenStudy (anonymous):

how could u find the velocity of the ruler right before the robot caught it

OpenStudy (anonymous):

Sure thang!

OpenStudy (anonymous):

would u use vf^2=vi^2+2ad?

OpenStudy (anonymous):

or how would u find it?

OpenStudy (anonymous):

Yeah, that's the equation you would use.

OpenStudy (anonymous):

ok would vi again be 0?

OpenStudy (anonymous):

and u would use the distance u got form the last question correct?

OpenStudy (anonymous):

is that right ? :D

OpenStudy (anonymous):

hello ? :o

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

s = 1/2 g t^2 should suffice

OpenStudy (amistre64):

the velocity? v = -gt is the usual setup

OpenStudy (amistre64):

convert cm to m, or m to cm along the way tho

OpenStudy (anonymous):

so it would be the same as using vf^2=vi^2+2ad wouldnt it?

OpenStudy (amistre64):

maybe .. if you want to go that route

OpenStudy (amistre64):

i dont have the physics formulas in memory, i just know it from a more mathical perspective

OpenStudy (anonymous):

would the answer be 39.2?

OpenStudy (anonymous):

cm/s squared

OpenStudy (amistre64):

for velocity? -9.8(.04) in m/s

OpenStudy (amistre64):

1m= 100cm ... so divide by 100 to get cm/s

OpenStudy (amistre64):

multiply not divide :)

OpenStudy (anonymous):

cant u just say its -981 isntead of 9.81

OpenStudy (amistre64):

i recall gravity being -9.8 m/s^2

OpenStudy (amistre64):

use whatever value for gravity your comfortable with

OpenStudy (amistre64):

m*cm/m = cm so multiply by 100 is the conversion factor

OpenStudy (anonymous):

yea so the velocity would be 39.2 cm/s^2 correct ? :D

OpenStudy (amistre64):

velocity is direction and magnitude, so -39.2 might be more proper

OpenStudy (anonymous):

yea i considered down to be positive for this lab so i didnt have to deal with negative numbers, but thank u guys so much!! :D